OpenStudy (samsan9):

For the system shown below, what are the coordinates of the solution that lies in quadrant I? x^2-y^2=25 x+y=25

OpenStudy (anonymous):

from second one, isolate x, what do you have?

OpenStudy (samsan9):

y=-x+25

OpenStudy (anonymous):

hihi, I suggest isolate x, you isolate y, but it's ok, now square both sides, show me what you have

OpenStudy (samsan9):

oh my bad its x=-y+2 and squaring both sides gives me x^2=Y^2+20

OpenStudy (anonymous):

no!!! the left hand side is x^2 but the right hand side is (-y+25)^2 or (25-y)^2 , open this perfect square, it is not y^2 +20, redo, please

OpenStudy (samsan9):

x^2=y^2-50y+100

OpenStudy (anonymous):

yup, now, take that x^2 replace to the first equation, what do you have?

OpenStudy (anonymous):